# references.bib

### v1

@@ Abstract
We consider the problem of learning the dynamics in the topology of time-evolving point clouds, the prevalent spatiotemporal model for systems exhibiting collective behavior, such as swarms of insects and birds or particles in physics. In such systems, patterns emerge from (local) interactions among self-propelled entities. While several well-understood governing equations for motion and interaction exist, they are difficult to fit to data due to the often large number of entities and missing correspondences between the observation times, which may also not be equidistant. To evade such confounding factors, we investigate collective behavior from a topological perspective, but instead of summarizing entire observation sequences (as in prior work), we propose learning a latent dynamical model from topological features per time point. The latter is then used to formulate a downstream regression task to predict the parametrization of some a priori specified governing equation. We implement this idea based on a latent ODE learned from vectorized (static) persistence diagrams and show that this modeling choice is justified by a combination of recent stability results for persistent homology. Various (ablation) experiments not only demonstrate the relevance of each individual model component, but provide compelling empirical evidence that our proposed model – neural persistence dynamics – substantially outperforms the state-of-the-art across a diverse set of parameter regression tasks.
@@ 1 Introduction
Understanding emerging behavioral patterns of a collective through the interaction of individual entities is key to elucidate many phenomena in nature on a macroscopic and microscopic scale. Prominent examples are coherently moving flocks of birds, the swarming behavior of insects and fish or the development of cancerous cells, all understood as 2D/3D point clouds that evolve over time. Importantly, several widely-accepted governing equations for collective behavior exist [Dorsogna06a, Vicsek95a, Motsch18a] which, when appropriately parameterized, can reproduce different incarnations of typically observed patterns. Even more importantly, these equations are tied to physically interpretable parameters and can provide detailed insights into the intrinsic mechanisms that control various behavioral regimes. However, while it is fairly straightforward to simulate collective behavior from governing equations (see [Diez21a]), the inverse problem, i.e., identifying the model parameters from the data, turns out to be inherently difficult. Confounding factors include the often large number of observed entities and the difficulty of reliably identifying individual trajectories due to missing correspondences between individuals at possibly non-equidistant observation times.
Figure 1: Neural Persistence Dynamics. Given is a sequence of observed point clouds $\mathcal{P}_{\tau_{0}},\ldots,\mathcal{P}_{\tau_{N}}$. First, we summarize each $\mathcal{P}_{\tau_{i}}$ via (Vietoris-Rips) persistent homology into persistence diagrams (zero-, one- and two-dimensional; only the zero-dimensional diagrams are shown here) which are then vectorized into $\mathbf{v}_{\tau_{i}}$ via existing techniques. Second, we model the dynamics in the sequence $\mathbf{v}_{\tau_{0}},\ldots,\mathbf{v}_{\tau_{N}}$ via a continuous latent variable model (in our case, a latent ODE) and then use the latent path to predict the parameters of well-established governing equation(s) of collective behavior. Precomputed steps are highlighted in red (best-viewed in color).
@@
Ours (PointNet++, v2) 0.272$\pm$0.083 0.199$\pm$0.014
Ours (PH, v1) 0.576$\pm$0.032 0.146$\pm$0.006
@@
Ours (PointNet++, v2) 0.814$\pm$0.032 0.132$\pm$0.018
Ours (PH, v1) 0.846$\pm$0.011 0.097$\pm$0.005
Yet, despite the often remarkable performance of topological approaches in terms of parameter identification for models of collective behavior, it remains unclear in which situations they are preferable over more traditional (learning) methods that can handle point cloud data, as, e.g., used in computer vision problems [Rempe20a, Guo20a]. In the spirit of [Turkes22a], we highlight this point by previewing a snapshot of one ablation experiment from Sec. 4. In particular, Tbl. 1 compares the parameter regression performance of our proposed approach, using PH, to a variant where we instead use (jointly learned) PointNet++ [Qi17a] representations. Referring to Fig. 1, this means that the $\mathbf{v}_{\tau_{i}}$ are computed by a PointNet++ model. As can be seen in Tbl. 1, relying on representations computed via PH works well on both datasets, while PointNet++ representations fail on one dataset (vicsek-10k), but indeed perform well on another (dorsogna-1k).
In view of this observation, it is worth highlighting, that there is a clear distinction in terms of the source of point cloud dynamics when comparing collective behavior to problems in computer vision where moving objects are of primary interest and PointNet variants (see [Guo20a]) have shown stellar performance: in particular, point clouds in vision primarily evolve due to a change in pose or camera motion, whereas collective behavior is driven by (local) intrinsic point interactions. The latter induces changes to the “shape” of the point clouds, i.e., a phenomenon that is typically not observed in vision.
@@ 2 Related work
Our work is partially related to the literature on learning with dynamic point clouds in vision problems, but primarily connects to work on summarizing topological features over time and inferring interaction laws of collective dynamics from data.
@@
Summarizing topological features over time. A common denominator in the literature on encoding topological changes in dynamic point clouds is the use of persistent homology, extended to accommodate the temporal dimension in various ways. One of the early works along this direction are persistence vineyards [CohenSteiner06a], introduced as a means to study folding trajectories of proteins. Based on the idea of tracking points in persistence diagrams over time, vineyards come with stability properties similar to persistent homology [Munch13a], but are expensive to compute and compare on a large scale. In cases where one would know the “right” scale at which to compute homology, one may also use zigzag persistence [Carlsson10a], as done in [Corcoran17a, Tymochko20a], to track homological changes over time. Yet, for problems with many observation sequences, scale selection per sequence is nontrivial and highly impractical.
@@
Alternatively, instead of thinking about the evolution of individual points in persistence diagrams one may discard the matching between points and instead consider sequences of summary representations of said diagrams. In [Hajij18a], the authors work directly with persistence diagrams to identify changes in the topology of time-varying graphs. In terms of temporal summary representations, [Topaz15a] introduce crocker plots to encode the evolution of topological features by stacking discretized Betti curves over time. Crocker stacks [Xian22a], an extension of this concept, adds a smoothing step that gradually reduces the impact of points of low persistence and, upon discretization, yields a third dimension to crocker plots. Both, crocker plots & stacks have been used as input to regression methods to predict the configuration of models of collective behavior [Bhaskar19a]. By drawing on prior work on kernels for sequentially ordered data [Kiraly19a][Giusti23a] follow a conceptually similar strategy as [Topaz15a, Xian22a], introducing a path signature kernel (PSK) for sequences of summary representations of persistence diagrams.
@@
Along a different line of research, [Kim18a, Kim18b] propose formigrams as summaries of dynamic metric data, encoding the evolution of connected components. In subsequent work, [Kim21a] construct multidimensional (i.e., spatiotemporal) persistent homology modules from dynamic metric data and compare invariants of these modules. While these works provide important theoretical stability results of dimension zero, i.e., connected components, it is unclear how their construction extends to homological features of higher dimension in a tractable manner. However, it is worth pointing out that recent progress for vectorizations of multiparameter persistent homology [Loiseaux2023a] in combination with [Kim21a] (who essentially construct multiparameter persistence modules) might, in the future, constitute a tractable approach to study dynamically changing point clouds with learning methods.
@@
Notably, computational challenges also arise in the context of crocker plots & stacks and PSKs. Despite remarkable success in distinguishing different configurations of models for collective behavior, all approaches suffer scalability issues, either in terms of the dimensionality of the vectorized persistence diagrams (as with the PSK approach of [Giusti23a]), or in terms of the number of observation sequences (as is the case for crocker plots/stacks, due to the need for extensive cross-validation of the discretization parameters). As we show in Sec. 4, our method outperforms these techniques by a large margin, scales to large amounts of training sequences and requires little hyperparameter tuning.
@@
Finally, aside from the closely-related works discussed above, we highlight that there is a large body of literature on the topological analysis of time-varying signals, such as studying fMRI data via cubical persistence [Rieck20a], or the persistent homology of delay embeddings [Perea15a, Perea15b] of time series. In our context, these works are only partially related, though, as they assume knowledge about individual trajectories over time (e.g., voxel IDs in fMRI data), which is unrealistic in our experimental setting.
@@
Inferring interaction laws for models of collective behavior. In the context of capturing the characteristics of collective behavior, there is a second line of closely related work on inferring and studying interaction laws (see, e.g., [Brunton16a, Ballerin08a, Bialek12a, Katz10a, Lu19a, Zhong20a]), ranging from metric-distance-based models and topological interaction models, to non-parametric estimators of interaction kernels. While a full survey of this literature is beyond the scope of this paper, we highlight that one common denominator in these works is their reliance on correspondences between points across time, e.g., to infer individual point velocities or trajectories. To give an example, in recent work [Zhong20a, Lu19a], the authors derive non-parametric estimators for interaction kernels (certain functions of pairwise distances) between observed points which crucially hinges on the traceability of each particle over time. While such approaches are conceptually appealing and even allow for reconstructing or continuing trajectories, they operate in the observation space. Even under a moderate number of points $m$, computing these estimators becomes prohibitively expensive (as, e.g., in 3D, the state space is $\mathbb{R}^{3m}$). In contrast, our approach only requires positional information, and can handle much larger point clouds. Furthermore, although we only present results on predicting parameters for a class of a priori given governing equations, by formulating an auxiliary regression task, our underlying model may also be used to predict other quantities of interest.
@@ 3 Method
Below, we present our method with a focus on its application to modeling collective behavior in point clouds. Note, however, that the latter represents only one of many conceivable use cases, as our method can be applied in exactly the same way, whenever one can extract topological features from time-dependent data (e.g., graphs or images).
@@
Problem statement. Given a sequence of point clouds $\mathcal{P}_{\tau_{0}},\ldots,\mathcal{P}_{\tau_{N}}$, observed at possibly non-equidistant time points $\tau_{i}$, we (1) seek to model their topological evolution over time and then (2) use this model to predict the parametrization of an a priori defined governing equation of collective behavior. The latter is typically specified by a small set of parameters $(\beta_{1},\ldots,\beta_{P})$ that control the motions $\mathrm{d}\mathbf{x}_{i}/\mathrm{d}t$ of individual points $\mathbf{x}_{i}$ and specify (local) interactions among neighboring points.
@@
As a preliminary introduction to our model description, we briefly establish how one may extract topological features from a point cloud $\mathcal{P}$, using persistent homology [Edelsbrunner2010Computational, Barannikov94a, Carlsson21a] – the arguably most prominent and computationally most feasible approach.
@@
Persistent homology of point clouds. Persistent homology seeks to uncover and concisely summarize topological features of $\mathcal{P}$. To this end, one constructs a topological space from $\mathcal{P}$ in the form of a simplicial complex, and studies its homology across multiple scales. The most relevant construction for our purposes is the Vietoris-Rips complex $\operatorname{Rips}(\mathcal{P})_{\delta}$, with vertex set $\mathcal{P}$. This complex includes an $m$-simplex $[\mathbf{x}_{0},\ldots,\mathbf{x}_{m}]$ iff $d(\mathbf{x}_{i},\mathbf{x}_{j})\leq\delta$ for all $0\leq i,j\leq m$ at a given threshold $\delta$. The “shape” of this complex can then be studied using homology, a tool from algebraic topology, with zero-dimensional homology ($H_{0}$) encoding information about connected components, one-dimensional homology ($H_{1}$) encoding information about loops and two-dimensional homology ($H_{2}$) encoding information about voids; we refer to this information as homological/topological features. Importantly, if $\delta_{b}\leq\delta_{d}$, then $\operatorname{Rips}(\mathcal{P})_{\delta_{b}}\subseteq\operatorname{Rips}(\mathcal{P})_{\delta_{d}}$, inducing a sequence of inclusions when varying $\delta$, called a filtration. The latter in turn induces a sequence of Abelian groups (or vector spaces) $\operatorname{H}_{k}(\operatorname{Rips}(\mathcal{P})_{\delta_{b}})\to\operatorname{H}_{k}(\operatorname{Rips}(\mathcal{P})_{\delta_{d}})$ at the homology level. Throughout this sequence, homological features (the generators of the groups) may appear and disappear, we say they are born at some $\delta_{b}$ and die at $\delta_{d}>\delta_{b}$. For instance, a one-dimensional hole might appear at a particular value of $\delta_{b}$ and disappear at a later $\delta_{d}$; in other words, it persists from $\delta_{b}$ to $\delta_{d}$, hence the name persistent homology. As different features may be born and die at the same time, the collection of (birth, death) tuples is a multiset of points, often represented in the form of a persistence diagram, which we denote as $\operatorname{dgm}_{k}(\operatorname{Rips}(\mathcal{P}))$. Here, the notation $\operatorname{Rips}(\mathcal{P})$ refers to the full filtration and $\operatorname{dgm}_{k}$ indicates that we have tracked $k$-dimensional homological features throughout this filtration.
@@
where $\mathbf{v}_{t_{i},k}\in\mathbb{R}^{d}$ and $d$ is controlled by hyperparameter(s) of the chosen vectorization technique. For brevity, we omit the subscript $k$ when referring to vectorization unless necessary.
@@
In case of vectorization techniques for persistence diagrams (e.g., [Adams17a, Hofer19b]), it turns out that, in most cases, vectorizations are only Lipschitz continuous with respect to Wasserstein distances, and existing results are typically of the form
@@
where $W_{1}$ denotes the $(1,2)$-Wasserstein distance as defined in [Skraba23a, Def. 2.7], $F,G$ are two persistence diagrams and $K>0$ is a Lipschitz constant. In the context of Rem. 1, one may be tempted to combine Eq. 2 with the seminal Wasserstein stability theorem from [CohenSteiner10a] to infer stability of vectorizations with respect to the input data. Yet, the conditions imposed in [CohenSteiner10a] actually eliminate a direct application of this result in most practical cases, as discussed in [Skraba23a]. However, the latter work also provides an alternative: in our particular case of Vietoris-Rips persistent homology, one can upper bound $W_{1}$ in terms of the standard point set Wasserstein distance $\mathcal{W}_{1}$. Specifically, upon relying on, e.g., the vectorization technique from [Hofer19b], we obtain the inequality chain
@@
where we set $\operatorname{vec}\equiv s_{\boldsymbol{\lambda}}$ as in [Hofer19b, Def. 9], and $F_{k}=\operatorname{dgm}_{k}(\operatorname{Rips}(\mathcal{P}))$, $G_{k}=\operatorname{dgm}_{k}(\operatorname{Rips}(\mathcal{Q}))$ denote the $k$-dimensional persistence diagrams of point clouds $\mathcal{P}$ and $\mathcal{Q}$ of equal cardinality $M$.
This continuity property guarantees that small changes in dynamic point clouds over time only induce small changes in their vectorized persistence diagrams, and therefore provides a solid justification for the model discussed next.
Latent variable model for persistence vectorizations. As, we presume that the dynamic point clouds under consideration are produced (or can be described sufficiently well) by equations of motions with only a few parameters, cf. Fig. 3, it is reasonable to assume that the dynamics of the (vectorized) persistence diagrams are equally governed by a somewhat simpler unobserved/latent dynamic process in $\mathbb{R}^{z}$ with $z\ll d$. We model this latent dynamic process via a neural ODE [Chen18a, Rubanova19a], learned in a variational Bayes regime [^1] Other choices, e.g., a latent SDE as in [Li20a, Zeng23a] are possible, but we did not explore this here. [Kingma14a]. In this setting, one chooses a recognition/encoder network (Encθ) to parametrize an approximate variational posterior $q_{\boldsymbol{\theta}}(\mathbf{z}_{t_{0}}|\{\mathbf{v}_{\tau_{i}}\}_{i})$, an ODE solver to yield $\{\mathbf{z}_{\tau_{i}}\}_{i}$ at observed time points $\tau_{i}\in[0,T]$ and a suitable generative/decoder network (Decγ) to implement the likelihood $p_{\boldsymbol{\gamma}}(\mathbf{v}_{\tau_{i}}|\mathbf{z}_{\tau_{i}})$. Upon choosing a suitable prior $p(\mathbf{z}_{t_{0}})$, one can then train the model via ELBO maximization, i.e.,
@@
Different to [Rubanova19a], we do not implement the recognition/encoder network via another neural ODE, but rather choose an attention-based approach (mTAN) [Shukla21a] which has proven to be very effective in practice [Zeng23a], and can even be used in a standalone manner as a strong baseline (see Sec. 4). In our implementation, the recognition network yields the parametrization ($\boldsymbol{\mu},\boldsymbol{\Sigma}$) of a multivariate Gaussian in $\mathbb{R}^{z}$ with diagonal covariance, and the prior is a standard Gaussian $\mathcal{N}(\mathbf{0},\mathbf{I}_{z})$. Furthermore, the ODE solver (e.g., Euler) can yield $\mathbf{z}_{t_{i}}$ at any desired $t_{i}$, however, we can only evaluate the ELBO at observed $\tau_{i}$.

### v2

# Neural Persistence Dynamics

@@ Abstract
We consider the problem of learning the dynamics in the topology of time-evolving point clouds, the prevalent spatiotemporal model for systems exhibiting collective behavior, such as swarms of insects and birds or particles in physics. In such systems, patterns emerge from (local) interactions among self-propelled entities. While several well-understood governing equations for motion and interaction exist, they are notoriously difficult to fit to data, as most prior work requires knowledge about individual motion trajectories, i.e., a requirement that is challenging to satisfy with an increasing number of entities. To evade such confounding factors, we investigate collective behavior from a topological perspective, but instead of summarizing entire observation sequences (as done previously), we propose learning a latent dynamical model from topological features per time point. The latter is then used to formulate a downstream regression task to predict the parametrization of some a priori specified governing equation. We implement this idea based on a latent ODE learned from vectorized (static) persistence diagrams and show that a combination of recent stability results for persistent homology justifies this modeling choice. Various (ablation) experiments not only demonstrate the relevance of each model component but provide compelling empirical evidence that our proposed model – Neural Persistence Dynamics – substantially outperforms the state-of-the-art across a diverse set of parameter regression tasks.
@@ 1 Introduction
Understanding emerging behavioral patterns of a collective through the interaction of individual entities is key to elucidate many phenomena in nature on a macroscopic and microscopic scale. Prominent examples are coherently moving flocks of birds, the swarming behavior of insects and fish or the development of cancerous cells, all understood as 2D/3D point clouds that evolve over time. Importantly, several widely-accepted governing equations for collective behavior exist [Dorsogna06a, Vicsek95a, Motsch18a] which, when appropriately parameterized, can reproduce different incarnations of typically observed patterns. Even more importantly, these equations are tied to physically interpretable parameters and can provide detailed insights into the intrinsic mechanisms that control various behavioral regimes. However, while it is fairly straightforward to simulate collective behavior from governing equations (see [Diez21a]), the inverse problem, i.e., identifying the model parameters from the data, turns out to be inherently difficult. Confounding factors include the often large number of observed entities and the difficulty of reliably identifying individual trajectories across point clouds at possibly non-equidistant observation times.
@@
Ours (PointNet++, v2) 0.274$\pm$0.085 0.199$\pm$0.014
Ours (PH, v1) 0.579$\pm$0.034 0.146$\pm$0.006
@@
Ours (PointNet++, v2) 0.816$\pm$0.031 0.132$\pm$0.018
Ours (PH, v1) 0.851$\pm$0.008 0.096$\pm$0.004
Yet, despite the often remarkable performance of topological approaches in terms of parameter identification for models of collective behavior, it remains unclear in which situations they are preferable over more traditional (learning) methods that can handle point cloud data, as, e.g., used in computer vision problems [Rempe20a, Guo20a]. In the spirit of [Turkes22a], we highlight this point by previewing a snapshot of one ablation experiment from Sec. 4. In particular, Tbl. 1 compares the parameter regression performance of our proposed approach, using PH, to a variant where we instead use PointNet++ [Qi17a] representations. Referring to Fig. 1, this means that the $\mathbf{v}_{\tau_{i}}$ are computed by a PointNet++ model. As can be seen in Tbl. 1, relying on representations computed via PH works well on both datasets, while PointNet++ representations fail on one dataset (vicsek-10k), but indeed perform reasonably well on the other (dorsogna-1k).
Figure 1: Conceptual overview of Neural Persistence Dynamics. Given is a sequence of observed point clouds $\mathcal{P}_{\tau_{0}},\ldots,\mathcal{P}_{\tau_{N}}$. First, we summarize each $\mathcal{P}_{\tau_{i}}$ via (Vietoris-Rips) persistent homology into persistence diagrams (zero-, one- and two-dimensional; only the zero-dimensional diagrams are shown) which are then vectorized into $\mathbf{v}_{\tau_{i}}$ via existing techniques. Second, we model the dynamics in the sequence $\mathbf{v}_{\tau_{0}},\ldots,\mathbf{v}_{\tau_{N}}$ via a continuous latent variable model (in our case, a latent ODE) and then use a summary of the latent path to predict the parameters of specific governing equation(s) of collective behavior. Precomputed steps are highlighted in red.
In light of this observation, it is worth emphasizing that there is a clear distinction in terms of the source of point cloud dynamics when comparing collective behavior to problems in computer vision where moving objects are of primary interest and PointNet variants (see [Guo20a]) have shown stellar performance: in particular, point clouds in vision primarily evolve due to a change in pose or camera motion, whereas collective behavior is driven by (local) intrinsic point interactions. The latter induces changes to the “shape” of the point clouds, i.e., a phenomenon that is typically not observed in vision.
@@ 2 Related work
Our work is partially related to the literature on learning with dynamic point clouds in vision problems, but primarily connects to work on summarizing topological features over time and inferring interaction laws of collective behavior from data.
@@
Summarizing topological features over time. A common denominator in the literature on encoding topological changes in dynamic point clouds is the use of persistent homology, extended to accommodate the temporal dimension in various ways. One of the early works along this direction are persistence vineyards [CohenSteiner06a], introduced as a means to study folding trajectories of proteins. Based on the idea of tracking points in persistence diagrams over time, vineyards come with stability properties similar to persistent homology [Munch13a], but are expensive to compute and compare on a large scale. In cases where one would know the “right” scale at which to compute homology, one may also use zigzag persistence [Carlsson10a], as done in [Corcoran17a, Tymochko20a], to track homological changes over time. Nevertheless, for problems with a large number of observation sequences, scale selection per sequence is nontrivial and highly impractical.
@@
Alternatively, instead of thinking about the evolution of individual points in persistence diagrams over time, one may discard the matching between points and instead focus on sequences of summary representations of said diagrams. In [Hajij18a], for instance, the authors work directly with persistence diagrams (per time point) to identify changes in the topology of time-varying graphs. In terms of temporal summary representations, [Topaz15a] introduce crocker plots to encode the evolution of topological features by stacking discretized Betti curves over time. Crocker stacks [Xian22a], an extension of this concept, adds a smoothing step that gradually reduces the impact of points of low persistence and, upon discretization, yields a third dimension to crocker plots. In our context, both crocker plots & stacks have been successfully used as input to regression methods to estimate the parametrization of models of collective behavior [Bhaskar19a]. By drawing on prior work on kernels for sequentially ordered data [Kiraly19a][Giusti23a] follow a conceptually similar strategy as [Topaz15a, Xian22a], introducing a path signature kernel (PSK) for sequences of summary representations of persistence diagrams.
@@
Along a different line of research, [Kim18a, Kim18b] propose formigrams as summaries of dynamic metric data, encoding the evolution of connected components. In subsequent work, [Kim21a] construct multidimensional (i.e., spatiotemporal) persistent homology modules from dynamic metric data and compare invariants of these modules. While these works provide important theoretical stability results for zero-dimensional homological features, i.e., connected components, it is unclear how their construction extends to homological features of higher dimension in a tractable manner. However, it is worth pointing out that recent progress along the lines of vectorizations of multiparameter persistent homology [Loiseaux2023a] in combination with [Kim21a] (who essentially construct multiparameter persistence modules) might, in the future, constitute a tractable approach to study dynamically changing point clouds with learning methods.
@@
Notably, computational challenges also arise in the context of crocker plots/stacks and PSKs. Despite their remarkable performance in distinguishing different configurations of models for collective behavior, both approaches suffer scalability issues: either (i) in terms of unfavorable scalability with respect to the dimensionality of vectorized persistence diagrams (as with the PSK approach of [Giusti23a]), or (ii) in terms of unfavorable scalability with the number of observation sequences (as is the case for crocker plots/stacks, due to the need for extensive cross-validation of the discretization parameters). As we show in Sec. 4, our method not only outperforms these techniques by a large margin, but also scales to large amounts of training sequences and requires little hyperparameter tuning.
@@
In addition to the closely related works discussed above, we highlight that there is a large body of literature on the topological analysis of time-varying signals, such as studying fMRI data via cubical persistence [Rieck20a], or the persistent homology of delay embeddings [Perea15a, Perea15b] of time series. In our context, however, these works are only partially related/relevant, as they do assume precise knowledge about individual trajectories over time (e.g., voxel IDs in fMRI data), which is unrealistic when seeking to infer parametrizations of models for collective behavior from data.
@@
Inferring interaction laws for models of collective behavior. In the context of studying characteristics of collective behavior, there is a second line of closely related work on inferring interaction laws (see, e.g., [Brunton16a, Ballerin08a, Bialek12a, Katz10a, Lu19a, Zhong20a]), ranging from metric-distance-based models and topological interaction models to non-parametric estimators of interaction kernels. While a thorough survey of this literature is beyond the scope of this paper, we highlight that one common denominator in these works is their reliance on correspondences between points across time, e.g., to infer individual point velocities or trajectories. To give an example, in recent work [Zhong20a, Lu19a], the authors derive non-parametric estimators for interaction kernels (certain functions of pairwise distances) between observed points which crucially hinges on the traceability of each particle over time. While such approaches are conceptually appealing and even allow for reconstructing or extrapolating trajectories, they operate in the observation space. Even under a moderate number of points $m$, computing these estimators becomes prohibitively expensive (as, e.g., in 3D, the state space is $\mathbb{R}^{3m}$). In contrast, our approach only requires positional information and can handle larger point clouds. Furthermore, although we only present results on predicting parameters for a class of a priori specified governing equations, by formulating an auxiliary regression task, our underlying model may also be used to predict other quantities of interest.
Finally, taking a slightly broader perspective on prior art, we want to highlight recent progress on learning-based approaches that study inverse problems in the context of classic partial differential equations (PDEs). It might be possible to apply such approaches [Zhao22a, Vadeboncoeur23a, Long24a] to specific models of collective behavior such as volume exclusion [Motsch18a], for which the asymptotics of infinitely many particles are solutions to parametric PDEs [Oelschlaeger90a]. However, for other models, the number of particles strongly influences the dynamics. For instance, in the D’Orsogna model [Dorsogna06a], particle distances can collapse to zero as the number of particles tends to infinity.
@@ 3 Method
Below, we present our method with a focus on its application to modeling collective behavior in point clouds. Importantly, the latter represents only one of many conceivable use cases. In fact, the core ideas can be applied in exactly the same way whenever one can extract topological features from time-dependent data (e.g., graphs or images).
@@
Problem statement. Given a sequence of point clouds $\mathcal{P}_{\tau_{0}},\ldots,\mathcal{P}_{\tau_{N}}$, observed at possibly non-equidistant time points $\tau_{i}$, we (1) seek to model their topological evolution over time and then (2) use this model to predict the parametrization of an a priori defined governing equation of collective behavior. The latter is typically specified by a small number of parameters $\beta_{1},\ldots,\beta_{P}$ that control the motions $\mathrm{d}\mathbf{x}_{i}/\mathrm{d}t$ of individual points $\mathbf{x}_{i}$ and specify (local) interactions among neighboring points.
@@
As preparation for our model description, we first briefly establish how one may extract topological features from a point cloud $\mathcal{P}$, using persistent homology [Edelsbrunner2010Computational, Barannikov94a, Carlsson21a] – the arguably most prominent and computationally most feasible approach.
@@
Persistent homology of point clouds. Persistent homology seeks to uncover and concisely summarize topological features of $\mathcal{P}$. To this end, one constructs a topological space from $\mathcal{P}$ in the form of a simplicial complex, and studies its homology across multiple scales. The most relevant construction for our purposes is the Vietoris-Rips complex $\operatorname{Rips}(\mathcal{P})_{\delta}$, with vertex set $\mathcal{P}$. This complex includes an $m$-simplex $[\mathbf{x}_{0},\ldots,\mathbf{x}_{m}]$ iff $d(\mathbf{x}_{i},\mathbf{x}_{j})\leq\delta$ for all $0\leq i,j\leq m$ at a given threshold $\delta$. The “shape” of this complex can then be studied using homology, a tool from algebraic topology, with zero-dimensional homology ($\operatorname{H}_{0}$) encoding information about connected components, one-dimensional homology ($\operatorname{H}_{1}$) encoding information about loops and two-dimensional homology ($\operatorname{H}_{2}$) encoding information about voids; we refer to this information as homological/topological features. Importantly, if $\delta_{b}\leq\delta_{d}$, then $\operatorname{Rips}(\mathcal{P})_{\delta_{b}}\subseteq\operatorname{Rips}(\mathcal{P})_{\delta_{d}}$, inducing a sequence of inclusions when varying $\delta$, called a filtration. The latter in turn induces a sequence of vector spaces $\operatorname{H}_{k}(\operatorname{Rips}(\mathcal{P})_{\delta_{b}})\to\operatorname{H}_{k}(\operatorname{Rips}(\mathcal{P})_{\delta_{d}})$ at the homology level. Throughout this sequence, homological features (corresponding to basis vectors of the different vector spaces) may appear and disappear; we say they are born at some $\delta_{b}$ and die at $\delta_{d}>\delta_{b}$. For instance, a one-dimensional hole might appear at a particular value of $\delta_{b}$ and disappear at a later $\delta_{d}$; in other words, the hole persists from $\delta_{b}$ to $\delta_{d}$, hence the name persistent homology. As different features may be born and die at the same time, the collection of (birth, death) tuples is a multiset of points, often represented in the form of a persistence diagram, which we denote as $\operatorname{dgm}_{k}(\operatorname{Rips}(\mathcal{P}))$. Here, the notation $\operatorname{Rips}(\mathcal{P})$ refers to the full filtration and $\operatorname{dgm}_{k}$ indicates that we have tracked $k$-dimensional homological features throughout this filtration.
@@
where $\mathbf{v}_{t_{i},k}\in\mathbb{R}^{d}$ and $d$ is controlled by hyperparameter(s) of the chosen vectorization technique. For brevity, we omit the subscript $k$ when referring to vectorizations unless necessary.
@@
In the case of vectorization techniques for persistence diagrams (e.g., [Adams17a, Hofer19b]), it turns out that, most of the time, vectorizations are only Lipschitz continuous with respect to Wasserstein distances, and existing results are typically of the form
@@
where $W_{1}$ denotes the $(1,2)$-Wasserstein distance as defined in [Skraba23a, Def. 2.7], $F,G$ are two persistence diagrams and $K>0$ is a Lipschitz constant. In the context of Rem. 1, one may be tempted to combine Eq. 2 with the seminal Wasserstein stability theorem from [CohenSteiner10a] to infer stability of vectorizations with respect to the input data. Yet, the conditions imposed in [CohenSteiner10a] actually eliminate a direct application of this result in most practical cases, as discussed in [Skraba23a]. However, the latter work also provides an alternative: in our particular case of Vietoris-Rips persistent homology, one can upper bound $W_{1}$ in terms of the standard point set Wasserstein distance $\mathcal{W}_{1}$. Specifically, upon relying on a stable vectorization technique, we obtain the inequality chain
@@
where $F_{k}=\operatorname{dgm}_{k}(\operatorname{Rips}(\mathcal{P}))$, $G_{k}=\operatorname{dgm}_{k}(\operatorname{Rips}(\mathcal{Q}))$ denote the $k$-dimensional persistence diagrams of point clouds $\mathcal{P}$ and $\mathcal{Q}$ of equal cardinality $M$. In particular, we realize $\operatorname{vec}$ via the approach outlined in [Hofer19b, Def. 10] using exponential structure elements [Hofer19b, Def. 19].
Overall, the continuity property in Eq. 3 guarantees that small changes in dynamic point clouds over time only induce small changes in their vectorized persistence diagrams, and therefore provides a solid justification for the model discussed next.
Latent variable model for persistence diagram vectorizations. As we presume that the dynamic point clouds under consideration are produced (or can be described sufficiently well) by equations of motions with only a few parameters, cf. Fig. 3, it is reasonable to assume that the dynamics of the (vectorized) persistence diagrams are equally governed by a somewhat simpler unobserved/latent dynamic process in $\mathbb{R}^{z}$ with $z\ll d$. We model this latent dynamic process via a neural ODE [Chen18a, Rubanova19a], learned in a variational Bayes regime [^1] Other choices, e.g., a latent SDE as in [Li20a, Zeng23a] are possible, but we did not explore this here. [Kingma14a]. In this setting, one chooses a recognition/encoder network (Encθ) to parametrize an approximate variational posterior $q_{\boldsymbol{\theta}}(\mathbf{z}_{t_{0}}|\{\mathbf{v}_{\tau_{i}}\}_{i})$, an ODE solver to yield latent states $\{\mathbf{z}_{\tau_{i}}\}_{i}$ at observed time points $\tau_{i}\in[0,T]$ and a suitable generative/decoder network (Decγ) to implement the likelihood $p_{\boldsymbol{\gamma}}(\mathbf{v}_{\tau_{i}}|\mathbf{z}_{\tau_{i}})$. Upon choosing a suitable prior $p(\mathbf{z}_{t_{0}})$, one can then train the model via ELBO maximization, i.e.,
@@
Different to [Rubanova19a], we do not implement the recognition/encoder network via another neural ODE, but rather choose an attention-based approach (mTAN) [Shukla21a] which can even be used in a standalone manner as a strong baseline (see Sec. 4). In our implementation, the recognition network yields the parametrization ($\boldsymbol{\mu},\boldsymbol{\Sigma}$) of a multivariate Gaussian in $\mathbb{R}^{z}$ with diagonal covariance, and the prior is a standard Gaussian $\mathcal{N}(\mathbf{0},\mathbf{I}_{z})$. Furthermore, the ODE solver (e.g., Euler) can yield $\mathbf{z}_{t_{i}}$ at any desired $t_{i}$, however, we can only evaluate the ELBO at observed time points $\tau_{i}$.